x^2y''-7xy'-20y=0 if y1=x^10
1
Expert's answer
2021-08-04T11:16:24-0400

Solution

Once we have this first solution we will then assume that a second solution will have the form

y2(x) = v(x)y1(x) or y2(x) = v(x)x10    

Differentiating

y2‘(x) = v’(x)x10 + 10v(x)x9 , y2‘’(x) = v’’(x)x10 + 20v’(x)x9 + 90v(x)x8    

Plugging these into the differential equation gives

v’’(x)x12 + 20v’(x)x11 + 90v(x)x10 - 7v’(x)x11 - 70v(x)x10 - 20v(x)x10  = 0   

v’’(x)x + 13v’(x) = 0   

Change of variable w(x) = v’(x), v’’(x) = w’(x)  =>  xw’ + 13 w = 0  => w(x) = Cx-13  =>  v(x) = -Cx-12/12+ D

Therefore   for C = -12 , D = 0 we obtain y2(x) = v(x)x10   = x-2   

Then general solution will then be

y(x) = Ax10 + Bx-2